I need a little help with this one. Thanks so much. The question asks to simplify:
(-2a^2b^3)^2 / (2a^2b)^3
i came up with ab^7 but i don't know for sure if it's correct. Any help would be appreciated.
: I need a little help with this one. Thanks so much. The question asks to simplify:
(-2a^2b^3)^2 / (2a^2b)^3
i came up with ab^7 but i don't know for sure if it's correct. Any help would be appreciated.

Answer by ankor@dixie-net.com(4668)

(-2a^2b^3)^2/(2a^2b)^3
:
I'm not sure how you got that answer, but let's just go thru it.
:
You multiply the numerator exponents by 2; (-2^2 = +4)
Multiply the denominator exponents by 3 (2^3=8) and we have
(4a^4b^6)/(8a^6b^3)
:
Combine the exponents of like terms, in a manner to ensure they are positive:
(b^((6-3)))/(2a^((6-4))) = b^3/(2a^2)
:
did this make sense to you?